OpenStudy (anonymous):

Imagine a basketball player is practicing free throw shots. For every 10 shots they make 70% of the time they make 8, 20% of the time they make 9, and 10% of the time they make all 10. For every 10 free throw shots they take, what is the expected number they will make?

OpenStudy (anonymous):

the number they make times the probability they make it

OpenStudy (anonymous):

\[.7\times 8+.2\times 9+.1\times 10\] in other words